# BillBoard Splatting (BBSplat): Learnable Textured Primitives for Novel View Synthesis

### v2

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. Our method’s qualitative and quantitative improvements over 3D and 2D Gaussians are most noticeable when fewer primitives are used, when BBSplat achieves over 1200 FPS. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. We demonstrate improvements on PSNR, SSIM, and LPIPS metrics compared to the state-of-the-art, especially for the case when fewer primitives are used, which, on the other hand, leads to up to 2×2$\times$2 × inference speed improvement for the same rendering quality. Project page: david-svitov.github.io/BBSplat_project_page.
@@ 1 Introduction
Recently, 2D Gaussian Splatting (2DGS) [19] proposed the use of flat Gaussians, oriented in 3D, to represent the scene more efficiently. Since 2D Gaussians are effectively tangent to object surfaces, they allow more accurate surface extraction. Despite proving their efficiency in mesh extraction task, 2D primitives result in a downgrade of rendering metrics compared to 3D primitives such as 3D Gaussians. In this work, we aim to make 2D primitives suitable for high-quality NVS by intoducing a new primitive representation.
Our proposed geometric primitives for NVS take inspiration from the classic billboards used for extreme 3D model simplification [9] by replacing mesh during a greedy optimization process. A 3D scene can be efficiently rendered by using a “billboard cloud” of given textured planar primitives with alpha channels. Using billboards, we can efficiently model planar surfaces, such as a painting on a wall or scene background, while dramatically reducing the number of geometric primitives required, up to an order of magnitude with respect to 3DGS/2DGS (check LABEL:fig:shapes for a comparison). In combination with efficient texture sampling implemented on the GPU, this leads to inference speed improvement without a drop in the rendering quality.
@@
To ensure rendering efficiency, we implemented texture sampling and a back-propagation process in CUDA. In this way, the use of textures does not result in overtime compared to the 2DGS Gaussians’ rasterization process. To tackle the challenge of storing all billboard textures, we compress them by representing each texture as sparse offsets from colors calculated by spherical harmonics [2] and by further quantizing them to 8 bits. Then, we can efficiently utilize dictionary-based compression algorithms [41, 11] for quantized textures.
@@
  • We propose BBSplat with optimizable textured primitives to learn 3D scene representation with photometric losses. BBSplat allows up to ×$\times 2$ acceleration for NVS compared to the state-of-the-art for the same rendering quality.
@@
  • We implemented BBSplat with CUDA and demonstrated its efficiency in extensive experiments on several real-scene open datasets.
@@ 2 Related work
Some of the recent works also propose to disentangle appearance and geometry for 3DGS. Texture-GS [49] introduces an MLP to establish correspondence between 3D Gaussian surface and texture map. In practice this approach allows one to edit the texture in the scene but decrease rendering speed to three times. Another approach [20] encodes color and opacity as spherical harmonics on the surface of 3D Gaussian, which allows one to improve reconstruction fidelity.
2D Gaussian splatting (2DGS) [19] was proposed to increase the quality of mesh surface reconstruction from such scene representation. Specifically, 2DGS uses ellipses instead of ellipsoids as they align along the object’s surface and make mesh extraction operation more straightforward. For faster rendering of such flat primitives, 2DGS utilizes a ray-splat intersection algorithm [48, 42]. While 2DGS increases the accuracy of mesh reconstruction, its fidelity falls shortly behind 3DGS, especially for in-the-wild scenes. In this work, we investigate the ability of flat primitives to be used for high-quality novel-view rendering.
We noted that existing approaches use a tremendous number of Gaussians to represent the scene, which affects rendering speed. Billboard clouds [9] propose an extreme 3D model simplification by replacing it with a set of textured planes. This representation is widely used in computer graphics for real-time forest rendering [3, 14, 27] and found implementation in such tasks as automatic paper slice-form design [28, 33, 39] and level of detail control [17]. In this work, we take inspiration from the billboard clouds technique [9] and propose arbitrarily shaped textured primitives as an extension of the 2DGS approach, resulting in faster and more qualitative rendering of real scenes reconstructed from the set of photographs.
@@ 3.5 Texture compression
Figure 3: Tanks&Temples metrics. Top row: We report quantitative metrics values depending on the number of primitives for each method. BBSplat outperforms state-of-the-art for the lower number of primitives. Bottom row: We demonstrate that lower number of primitives (indicated by marker size) leads to rendering speed (frames per second - FPS) improvement. BBSplat demonstrates a better metrics/speed ratio for the NVS task (better results direction shown by the grey arrow).
@@ 4 Experiments
Table 2: Quantitative results with a large number of primitives. We report metrics for the following datasets: Tanks&Temples [26], Mip-NeRF-360 [5], and DTU [22]. We provide metrics for the recommended number of Gaussians for each method, which results in lower FPS and increased storage space.†- indicates method version without using depth-normal regularizations.
@@
Texture-GS 30.03 0.0135 0.1440 267
Ours-small 30.81 0.0123 0.1055 1230
@@ 4.2 Results and Evaluation
Quantitative comparison. We report PSNR, SSIM, and LPIPS [55] metrics (Tabs. 1 and 2) widely used to evaluate quality in the NVS task. PSNR and SSIM are traditionally used to demonstrate similarity with ground truth images; LPIPS utilize neural networks and is known for the best correlation with human perception. To assess model efficiency, we also report rendering speed as frames per second (FPS) and required storage space along with a number of used primitives. In the provided experiments, we evaluated metrics for different numbers of primitives by limiting the growth of their number during training as proposed in [6].
@@
In an experiment in Table 1, we showcase metrics for the fixed number of primitives depending only on their type. For the same amount of primitives, our method reaches better PSNR, SSIM and LPIPS than state-of-the-art Gaussian-based representations both for indoor and outdoor scenes. While improving state-of-the-art objective metrics, for the same number of primitives BBSplat requires more storage space. In the next experiment, we demonstrate inference speed and storage efficiency for on-par metrics.
@@
In Table 2, we compare against other methods for the maximum recommended amount of Gaussians to achieve the best possible quality for each method. For our method, we provide two scenarios with fewer (small) and more (big) billboards. We outperform 3DGS and 2DGS for all three metrics, at the same time showing inference speed and storage space improvement. BBSplat demonstrates metrics on par with the best competitor, 3DGS-MCMC, achieving up to two times faster rendering of the scenes with three times lower storage space requirement.
@@
In Table 3, we compare with Texture-GS [49] as it also proposes to employ textures, although in conjunction with a 3DGS model. The main objective of Texture-GS is texture editing: indeed we verify that Texture-GS performance are reduced for the NVS task, especially in FPS count. We report metrics only for the DTU dataset since Texture-GS is unable to process large scenes, as it is limited by spherical texture space definition.
@@
In Figure 3, we show the metrics as a function of the number of primitives (Gaussians or BBSplats) and as a function of FPS. To plot the leftmost point we used the maximum number of Gaussians requested by each method, then for subsequent points we fixed the number of primitives for all scenes at the chosen value. It can be seen that our method gives a flatter curve of metrics change, which means a minimal decrease in metrics when the number of primitives decreases to 80K. Thus, the proposed implementation allows us to achieve an inference speed of up to 1250 FPS while maintaining high rendering quality.
@@
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art while providing faster inference speed and higher level of detail. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to Sec. 7.1 of supp. mat..
@@ 5 Conclusion
We have proposed BBSplat - a novel method for 3D scene representation. The proposed method uses textured primitives of arbitrary learnable shapes and allows a two times faster generation of photorealistic novel views. We developed specialized regularization term and compression technique to reduce storage space by leveraging the sparse nature of the proposed textured representation. In extensive experiments on real data, we demonstrated the efficiency of the proposed method both quantitatively and qualitatively. In particular, we showed the best trade-off between objective metrics and inference speed.
@@ 6.2 Hyper-parameters
Table 6 shows metrics for different texture sizes and number of billboards. We select texture size to be the power of two (8, 16, 32, 64) as it is more suitable for processing with GPU, and set the corresponding number of billboards to result in 40.96M parameters for one texture channel to not exceed GPU capacity. While $32\times 32$ and $16\times 16$ both provide suitable metrics, we chose $16\times 16$ as slightly more accurate. As an alternative, $32\times 32$ can be used to achieve faster inference.
@@ 7.1 Visual results
In Figure 7, we provide an additional visual comparison of BBSplat with state-of-the-art NVS methods. Our renderings demonstrate exceeding quality for similar or higher FPS.

### v3

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. The proposed primitives close the rendering quality gap between 2D and 3D Gaussian Splatting (GS), preserving the accurate mesh extraction ability of 2D primitives. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in the storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. Project page: david-svitov.github.io/BBSplat_project_page.
@@ 1 Introduction
Recently, 2D Gaussian Splatting (2DGS) [20] proposed the use of flat Gaussians, oriented in 3D as flat primitives align better with the object surface. Since 2D Gaussians are effectively tangent to object surfaces, they allow more accurate surface extraction. Despite proving their efficiency in mesh extraction task, 2D primitives result in a downgrade of rendering metrics compared to 3D primitives such as 3D Gaussians. In this work, we aim to make 2D primitives suitable for high-quality NVS by intoducing a new primitive representation.
Our proposed geometric primitives for NVS take inspiration from the classic billboards used for extreme 3D model simplification [9] by replacing mesh during a greedy optimization process. A 3D scene can be efficiently rendered using a “billboard cloud” of textured planar primitives with alpha channels. Using billboards, we can efficiently model planar surfaces, such as a painting on a wall or scene background, while dramatically reducing the number of geometric primitives required, up to an order of magnitude with respect to 3DGS/2DGS (check LABEL:fig:shapes for a comparison).
@@
To tackle the challenge of storing all billboard textures, we compress them by representing each texture as sparse offsets from colors calculated by spherical harmonics [2] and by further quantizing them to 8 bits. Then, we can efficiently utilize dictionary-based compression algorithms [42, 11] for quantized textures.
@@
  • We propose BBSplat with optimizable textured primitives to learn 3D scene representation with photometric losses. The proposed representation surpasses 2DGS and 3DGS in rendering quality, especially for smaller numbers of primitives.
  • The BBSplat representation allows extracting accurate meshes from a scene with accuracy superior to 3D Gaussian-based methods.
@@ 2 Related work
Some of the recent works also propose to disentangle appearance and geometry for 3DGS. Texture-GS [50] introduces an MLP to establish correspondence between 3D Gaussian surface and texture map. In practice this approach allows one to edit the texture in the scene but is limited to synthetic scenes with one object and no background. Another approach [21] encodes color and opacity as spherical harmonics on the surface of 3D Gaussian, which allows one to improve reconstruction fidelity.
2D Gaussian splatting (2DGS) [20] was proposed to increase the quality of mesh surface reconstruction from such scene representation. Specifically, 2DGS uses ellipses instead of ellipsoids as they align along the object’s surface and make mesh extraction operation more straightforward. For rendering of such flat primitives, 2DGS utilizes a ray-splat intersection algorithm [49, 43]. While 2DGS increases the accuracy of mesh reconstruction, its fidelity falls shortly behind 3DGS, especially for in-the-wild scenes. In this work, we investigate the ability of flat primitives to be used for high-quality novel-view rendering.
Billboard clouds [9] propose an extreme 3D model simplification by replacing it with a set of textured planes. This representation is widely used in computer graphics for real-time forest rendering [3, 14, 28] and found implementation in such tasks as automatic paper slice-form design [29, 34, 40] and level of detail control [18]. In this work, we take inspiration from the billboard clouds technique [9] and propose arbitrarily shaped textured primitives as an extension of the 2DGS approach.
@@ 3.5 Texture compression
Figure 3: Tanks&Temples metrics. We report quantitative metrics values depending on the number of primitives for each method. BBSplat outperforms state-of-the-art for the lower number of primitives.
@@ 4 Experiments
Table 2: Quantitative results with a large number of primitives. We report metrics for the following datasets: Tanks&Temples [27], Mip-NeRF-360 [5], and DTU [23]. We provide metrics for the recommended number of Gaussians for each method, which results in increased storage space.†- indicates method version without using depth-normal regularizations.
@@
Texture-GS 30.03 0.0135 0.1440
Ours-small 30.81 0.0123 0.1055
@@ 4.2 Results and Evaluation
Quantitative comparison. We report PSNR, SSIM, and LPIPS [56] metrics (Tabs. 1 and 2) widely used to evaluate quality in the NVS task. PSNR and SSIM are traditionally used to demonstrate similarity with ground truth images; LPIPS utilize neural networks and is known for the best correlation with human perception. In the provided experiments, we evaluated metrics for different numbers of primitives by limiting the growth of their number during training as proposed in [6].
@@
In an experiment in Table 1, we showcase metrics for the fixed number of primitives depending only on their type. For the same amount of primitives, our method reaches better PSNR, SSIM and LPIPS than state-of-the-art Gaussian-based representations both for indoor and outdoor scenes. While improving state-of-the-art metrics, for the same number of primitives BBSplat requires more storage space.
@@
In Table 2, we compare against other methods for the maximum recommended amount of Gaussians to achieve the best possible quality for each method. For our method, we provide two scenarios with fewer (small) and more (big) billboards.
@@
In Table 3, we compare with Texture-GS [50] as it also proposes to employ textures, although in conjunction with a 3DGS model. We report metrics only for the DTU dataset since Texture-GS is unable to process large scenes, as it is limited by spherical texture space definition.
@@
In Figure 3, we show the metrics as a function of the number of primitives (Gaussians or BBSplats). To plot the leftmost point we used the maximum number of Gaussians requested by each method, then for subsequent points we fixed the number of primitives for all scenes at the chosen value. It can be seen that our method gives a flatter curve of metrics change, which means a minimal decrease in metrics when the number of primitives decreases to 80K.
@@
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to Sec. 7.1 of supp. mat..
Mesh Extraction. In Table 4, we compare the 3D mesh extraction performance of BBSPlat with state-of-the-art methods on the DTU dataset [23]. For the experiment, we use 15 scenes with ground truth obtained by 3D scanning objects in a laboratory setting. As a metric, we present the Chamfer distance indicating the similarity between two meshes. Both experiments for 2DGS and for BBSplat were conducted using the depth-normal regularization proposed in [20]. Rendering quality metrics for BBSplat in Tables 1 and 2 were also obtained using such regularizations. BBSplat reaches 0.91 Chamfer distance on DTU dataset, which is better than 1.96 for 3DGS and 1.33 for SuGaR, and close to 0.80 for 2DGS.
Method Chamfer distance
3DGS [24] 1.96
SuGaR [17] 1.33
2DGS [20] 0.80
BBSplat(Ours) 0.91
Table 4: Chamfer distance on the DTU dataset [23] We report the Chamfer distance on 15 scenes from DTU dataset to evaluate 3D mesh extraction quality of the method.
@@ 5 Conclusion
We have proposed BBSplat - a novel method for 3D scene representation. The proposed method uses textured primitives of arbitrary learnable shapes and allows accurate mesh extraction and photorealistic novel views synthesis of real scenes. We developed specialized regularization term and compression technique to reduce storage space by leveraging the sparse nature of the proposed textured representation. In extensive experiments on real data, we demonstrated the efficiency of the proposed method both quantitatively and qualitatively.
@@ 6.2 Hyper-parameters
Table 7 shows metrics for different texture sizes and number of billboards. We select texture size to be the power of two (8, 16, 32, 64) as it is more suitable for processing with GPU, and set the corresponding number of billboards to result in 40.96M parameters for one texture channel to not exceed GPU capacity. While $32\times 32$ and $16\times 16$ both provide suitable metrics, we chose $16\times 16$ as slightly more accurate.
@@ 7.1 Visual results
In Figure 7, we provide an additional visual comparison of BBSplat with state-of-the-art NVS methods. Our renderings demonstrate exceeding quality.